If 1.22 kJ of heat is added to 60.0 g of water at 29.0 ◦C, what is the final temperature of the water? (The specific heat of water is 4.184 J/g ◦C)

An unknown metal with a mass of 36.0 g absorbs 58 J of heat. Its temperature rises from 10.4 ◦C to 12.9 ◦C. What is the specific heat of the unknown metal?
1
Expert's answer
2020-09-14T08:28:43-0400

(1)"\; ∆H = mC(T_2-T_1)"

"m = 60 \;g"

"C = 4.184 \;J/g ◦C"

"T_1 = 29.0\; ◦C"

"∆H = 1220\; J"

"1220 = 60\times 4.184\times(T_2 – 29.0)"

"4.86 = T_2 \u2013 29.0"

"T_2 = 33.86 \;◦C"

(2) "\u2206H = mC(T_2-T_1)"

"m = 36.0 \;g"

"T_1 = 10.4\; ◦C"

"T_2 = 12.9\; ◦C"

"∆H = 58 \;J"

"58 = 36.0\times C\times (12.9-10.4)"

"58 = 90\;C"

"C = 0.64 \;J/g ◦C"